# Gaussian Mixture Models Based Augmentation Enhances GNN Generalization

### v2

@@ 5 Conclusion
$\frac{d\hat{\theta}_{\epsilon_{n,m}}}{d\epsilon_{n,m}}\simeq\frac{\hat{\theta}_{\epsilon_{n,m}}-\hat{\theta}}{\epsilon_{n,m}}=-\left[\nabla_{\hat{\theta}}^{2}\mathcal{L}(\hat{\theta})\right]^{-1}\nabla_{\hat{\theta}}\ell(\mathcal{G}_{n}^{\lambda_{n,m}},\hat{\theta}).$
$\frac{d\ell(\mathcal{G}_{k}^{test},\hat{\theta}_{\epsilon_{n,m}})}{d\epsilon_{n,m}}$ $=\frac{d\ell(\mathcal{G}_{k}^{test},\hat{\theta}_{\epsilon_{n,m}})}{d\hat{\theta}_{\epsilon_{n,m}}}\frac{d\hat{\theta}_{\epsilon_{n,m}}}{d\epsilon_{n,m}}.$
Appendix C Configuration Models
@@
In this section, we present a novel adaptation of Configuration Models as a graph data augmentation technique for GNN. Configuration Models Newman (2013) enable the generation of randomized graphs that maintain the original degree distribution. We can, therefore, leverage this strategy to improve the generalization of GNNs. Below, we present the steps involved in our approach to using Configuration Models for Graph Data Augmentation:
@@
    Stub Creation: Using a Bernoulli distribution with parameter $p\in[0,1]$, we randomly select a subset of candidate edges and break them to create stubs (half-edges).
@@
Table 4 shows the performance of this approach on the two GNN backbone GCN and GIN.
Table 4: Classification accuracy ($\pm$ standard deviation) on different benchmark node classification datasets for the data augmentation baselines based on the GIN backbone. The higher the accuracy (in %) the better the model.
Model IMDB-BINARY IMDB-MULTI MUTAG PROTEINS DD
Config Models w/ GCN 71.70 (3.16) 48.40 (3.88) 74.97 (6.77) 70.08 (4.93) 69.01 (3.44)
Config Models w/ GIN 71.70 (4.24) 49.00 (3.44) 81.43 (10.05) 68.34 (5.30) 71.61 (5.96)
Appendix D Ablation Study
To provide additional comparison and motivate the use of GMMs with the EM algorithm, we expanded our evaluation to include additional methods for modeling the distribution of the graph representations. Specifically, the comparison includes:
@@ Appendix D Ablation Study
Table 5: Ablation Study on the density estimation scheme for learned GCN representations.
Model IMDB-BINARY IMDB-MULTI MUTAG PROTEINS DD
GMM w/ EM 71.00 (4.40) 49.82 (4.26) 76.05 (6.47) 70.97 (5.07) 71.90 (2.81)
GMM w/ VBI 71.00 (4.21) 49.53 (4.26) 76.05 (6.47) 70.97 (4.52) 71.64 (2.90)
KDE 55.90 (10.29) 39.53 (2.87) 66.64 (6.79) 59.56 (2.62) 58.66 (3.97)
Copula 69.80 (4.04) 47.13 (3.45) 74.44 (6.26) 65.04 (3.37) 65.70 (3.04)
GAN 70.60 (3.41) 48.80 (5.51) 75.52 (4.96) 69.98 (5.46) 66.26 (3.72)
Table 6: Ablation Study on the density estimation scheme for learned GIN representations.
Model IMDB-BINARY IMDB-MULTI MUTAG PROTEINS DD
GMM w/ EM 71.70 (4.24) 49.20 (2.06) 88.83 (5.02) 71.33 (5.04) 68.61 (4.62)
GMM w/ VBI 71.40 (2.65) 47.80 (2.22) 88.30 (5.19) 70.25 (4.65) 67.82 (4.96)
KDE 69.10 (3.93) 41.46 (3.02) 77.60 (6.83) 60.37 (3.04) 67.48 (6.18)
Copula 70.60 (2.61) 47.60 (2.29) 88.30 (5.19) 70.16 (4.55) 67.91 (4.90)
GAN 70.50 (3.80) 48.40 (1.71) 88.83 (5.02) 71.33 (5.55) 67.74 (4.82)
@@
Appendix E Training and Augmentation time
We compare the data augmentation times of our approach and the baselines in Table 7. In addition to outperforming the baselines on most datasets, our approach offers an advantage in terms of time complexity. The training time of baseline models varies depending on the augmentation strategy used, specifically, whether it involves pairs or individual graphs. Even in cases where a graph augmentation has a low computational cost for some baselines, training can still be time-consuming as multiple augmented graphs are required to achieve satisfactory test accuracy. For instance, methods like DropEdge, DropNode, and SubMix, while computationally simple, require generating multiple augmented samples at each epoch, thereby increasing the overall training time. In contrast, GMM-GDA introduces a more efficient approach by generating only one augmented graph per training instance, which is reused across all epochs. This design ensures a balance between computational efficiency and augmentation effectiveness, reducing the overall training burden while maintaining strong performance. The only baseline that is more time-efficient than our approach is GeoMix; however, our method consistently outperforms GeoMix across all settings, as shown in Tables 1 and 2.
@@ Appendix E Training and Augmentation time
Attack Model IMDB-BINARY MUTAG DD

### v3

@@ Acknowledgment
$\frac{d\hat{\theta}_{\epsilon_{n,m}}}{d\epsilon_{n,m}}\simeq\frac{\hat{\theta}_{\epsilon_{n,m}}-\hat{\theta}}{\epsilon_{n,m}}=-\left[\nabla_{\hat{\theta}}^{2}\mathcal{L}(\hat{\theta})\right]^{-1}\nabla_{\hat{\theta}}\ell(\widetilde{\mathcal{G}}_{n}^{m},\hat{\theta}).$
$\frac{d\ell(\mathcal{G}_{k}^{test},\hat{\theta}_{\epsilon_{n,m}})}{d\epsilon_{n,m}}$ $=\frac{d\ell(\mathcal{G}_{k}^{test},\hat{\theta}_{\epsilon_{n,m}})}{d\hat{\theta}_{\epsilon_{n,m}}}\frac{d\hat{\theta}_{\epsilon_{n,m}}}{d\epsilon_{n,m}}$
$=-\nabla_{\theta}\ell(\mathcal{G}_{k}^{test},\hat{\theta})\mathbf{H}_{\hat{\theta}}^{-1}\nabla_{\theta}\ell(\widetilde{\mathcal{G}}_{n}^{m},\hat{\theta}).$
Appendix D Mathematical Expressions of GCN and GIN
In this section, we provide concise definitions of two widely used GNN architectures: Graph Convolutional Networks (GCN) and Graph Isomorphism Networks (GIN). These architectures differ in how they aggregate and combine information from neighboring nodes in a graph.
Graph Convolutional Network (GCN) (Kipf & Welling, 2017).
The GCN updates node embeddings by aggregating normalized features from their neighbors. Specifically, for a node $v\in\mathcal{V}$, its feature vector $\mathbf{h}_{v}^{(t)}$ at layer $t$ is computed as,
$\mathbf{h}_{v}^{(t)}=\sigma\Bigg{(}\sum_{u\in\mathcal{N}(v)\cup\{v\}}\frac{1}{\sqrt{\deg(v)\deg(u)}}\mathbf{W}^{(t)}\mathbf{h}_{u}^{(t-1)}\Bigg{)},$
where $\mathbf{h}_{u}^{(t-1)}$ is the feature vector of node $u$ at layer $t-1$, $\mathbf{W}^{(t)}$ is a trainable weight matrix for layer $t$, and $\sigma(\cdot)$ is a non-linear activation function, such as ReLU.
Graph Isomorphism Network (GIN) (Xu et al., 2019).
GIN is designed to match the expressiveness of the Weisfeiler-Lehman (WL) graph isomorphism test. It updates a node’s embedding by aggregating its own feature with those of its neighbors, followed by a a multi-layer perceptron (MLP). The update rule at each message passing layer $t$ is
$\mathbf{h}_{v}^{(t)}=\text{MLP}^{(t)}\Big{(}(1+\epsilon^{(t)})\cdot\mathbf{h}_{v}^{(t-1)}+\sum_{u\in\mathcal{N}(v)}\mathbf{h}_{u}^{(t-1)}\Big{)},$
where $\epsilon^{(t)}$ is a fixed or learnable scalar. GIN allows for more expressive feature transformations compared to methods with fixed aggregation schemes, enabling it to distinguish a broader range of graph structures.
Matrix forms and comparison.
The node representations $\mathbf{h}_{v}^{(t)}$ at each layer can be expressed in matrix form as $\mathbf{H}^{(t)}\in\mathbb{R}^{p\times d_{t}}$, where $p$ is the number of nodes in the graph, $d_{t}$ is hidden dimension in the $t-$th layer, and $\mathbf{H}^{(t)}$ is the concatenation of all node representations $\mathbf{h}_{v}^{(t)}$ for $v\in\mathcal{V}$. For GCN, the update rule can be written as:
$\mathbf{H}^{(t)}=\sigma\Bigl{(}\widetilde{\mathbf{D}}^{-1/2}\widetilde{\mathbf{A}}\widetilde{\mathbf{D}}^{-1/2}\mathbf{H}^{(t-1)}\mathbf{W}^{(t)}\Bigr{)},$
where $\widetilde{\mathbf{A}}=\mathbf{A}+\mathbf{I}$ is the adjacency matrix with self-loops, and $\widetilde{\mathbf{D}}$ is its diagonal degree matrix. In contrast, for GIN, the update rule is given by:
$\mathbf{H}^{(t)}=\mathrm{MLP}^{(t)}\Bigl{(}(1+\epsilon^{(t)})\mathbf{H}^{(t-1)}+\mathbf{A}\mathbf{H}^{(t-1)}\Bigr{)}.$
If the same MLP is used, comprising a learnable linear layer followed by a ReLU activation, the only difference between GCN and GIN lies in the graph shift operator: GCN uses the degree-normalized operator $\widetilde{\mathbf{D}}^{-1/2}\widetilde{\mathbf{A}}\widetilde{\mathbf{D}}^{-1/2}$, while GIN uses $(1+\epsilon^{(t)})\mathbf{I}+\mathbf{A}$.
Therefore, when node features are taken as constant and only the graph structure is considered, the difference in their Lipschitz behavior can be traced back to the function that maps the adjacency matrix $A$ (and degrees) to these respective shift operators. In other words, the Lipschitz constant difference arises from whether the adjacency information is normalized ($\widetilde{\mathbf{D}}^{-1/2}\,\widetilde{\mathbf{A}}\,\widetilde{\mathbf{D}}^{-1/2}$) or as ($(1+\epsilon^{(t)})\mathbf{I}+\mathbf{A}$).
Lipschitz constants of GCN and GIN.
Under constant node features, the Lipschitz constant of each architecture can be decomposed into two parts: one accounting for the sensitivity of the mapping from the adjacency matrix to the respective graph shift operator, i.e., degree-normalized $\widetilde{\mathbf{D}}^{-1/2}\,\widetilde{\mathbf{A}}\,\widetilde{\mathbf{D}}^{-1/2}$ for GCN vs. $(1+\epsilon)\mathbf{I}+\mathbf{A}$ for GIN, and the other capturing all shared learnable transformations.
Let $\ell_{\mathrm{GCN}}$ and $\ell_{\mathrm{GIN}}$ be respectively the graph shift operators $A\mapsto\widetilde{\mathbf{D}}^{-1/2}\,\widetilde{\mathbf{A}}\,\widetilde{\mathbf{D}}^{-1/2}$ and $A\mapsto(1+\epsilon)\mathbf{I}+\mathbf{A}$, and let $\ell_{\mathrm{params}}$ represent the product of Lipschitz factors arising from the shared functions. Then,
$L_{\mathrm{GCN}}$ $=\ell_{\mathrm{GCN}}\times\ell_{\mathrm{params}},$
$L_{\mathrm{GIN}}$ $=\ell_{\mathrm{GIN}}\times\ell_{\mathrm{params}}.$
It becomes clear that the difference between GCN and GIN Lipschitz constants depends solely on whether the graph adjacency is normalized, since $\ell_{\mathrm{params}}$ is common to both. It’s straightforward that $\ell_{\mathrm{GIN}}\leq 1$ since the corresponding graph shift operator is just a translation. Let us now derive an upper bound for $\ell_{\mathrm{GIN}}.$ Let $\mathbf{A}_{1},\mathbf{A}_{2}\in\mathbb{R}^{p\times p}$ be adjacency matrices of graphs on $n$ nodes. For each adjacency matrix $A_{i}$, we define the diagonal degree matrix, $\mathbf{D}_{i}=\mathrm{diag}\bigl{(}\text{deg}_{1}^{(i)},\dots,\text{deg}_{p}^{(i)}\bigr{)}$ where $\forall j\leq p,~{}~{}\text{deg}_{j}^{(i)}=\sum_{k=1}^{n}(\mathbf{A}_{i})_{jk}.$ We have,
$D_{1}^{-\tfrac{1}{2}}\,\mathbf{A}_{1}\,\mathbf{D}_{1}^{-\tfrac{1}{2}}-\mathbf{D}_{2}^{-\tfrac{1}{2}}\,\mathbf{A}_{2}\,\mathbf{D}_{2}^{-\tfrac{1}{2}}$ $=\mathbf{D}_{1}^{-\tfrac{1}{2}}\,\mathbf{A}_{1}\,\mathbf{D}_{1}^{-\tfrac{1}{2}}-\mathbf{D}_{1}^{-\tfrac{1}{2}}\,\mathbf{A}_{2}\,\mathbf{D}_{1}^{-\tfrac{1}{2}}+\mathbf{D}_{1}^{-\tfrac{1}{2}}\,\mathbf{A}_{2}\,\mathbf{D}_{1}^{-\tfrac{1}{2}}\;-\;\mathbf{D}_{2}^{-\tfrac{1}{2}}\,\mathbf{A}_{2}\,\mathbf{D}_{2}^{-\tfrac{1}{2}}$
$=\underbrace{\mathbf{D}_{1}^{-\tfrac{1}{2}}\,\bigl{(}\mathbf{A}_{1}-\mathbf{A}_{2}\bigr{)}\,\mathbf{D}_{1}^{-\tfrac{1}{2}}}_{\text{(i)}}+\underbrace{\bigl{[}\mathbf{D}_{1}^{-\tfrac{1}{2}}-\mathbf{D}_{2}^{-\tfrac{1}{2}}\bigr{]}\;\mathbf{A}_{2}\;\mathbf{D}_{1}^{-\tfrac{1}{2}}+\mathbf{D}_{2}^{-\tfrac{1}{2}}\,\mathbf{A}_{2}\,\bigl{[}\mathbf{D}_{1}^{-\tfrac{1}{2}}-\mathbf{D}_{2}^{-\tfrac{1}{2}}\bigr{]}}_{\text{(ii)}}.$
We can upperbound each of (i), and (ii). For (i), we have the following upperbound,
$\bigl{\|}\mathbf{D}_{1}^{-\tfrac{1}{2}}\,\bigl{(}\mathbf{A}_{1}-\mathbf{A}_{2}\bigr{)}\,\mathbf{D}_{1}^{-\tfrac{1}{2}}\bigr{\|}\leq\|\mathbf{D}_{1}^{-\tfrac{1}{2}}\|^{2}\,\|\mathbf{A}_{1}-\mathbf{A}_{2}\|\leq\frac{1}{\delta_{1,\text{min}}}\,\|\mathbf{A}_{1}-\mathbf{A}_{2}\|,$
[…]
$\|(ii)\|$ $\leq\|\mathbf{D}_{1}^{-1/2}-\mathbf{D}_{2}^{-1/2}\|\|\mathbf{A}_{2}\|\bigl{(}\|\mathbf{D}_{1}^{-1/2}\|+\|\mathbf{D}_{2}^{-1/2}\|\bigr{)},$
$\leq\frac{2}{\text{min}(\delta_{1,\text{min}},\delta_{2,\text{min}})^{1/2}}\|\mathbf{D}_{1}^{-1/2}-\mathbf{D}_{2}^{-1/2}\|\|\mathbf{A}_{2}\|$
$\leq\frac{2M}{\text{min}(\delta_{1,\text{min}},\delta_{2,\text{min}})^{1/2}}\|\mathbf{D}_{1}^{-1/2}-\mathbf{D}_{2}^{-1/2}\|$
$\leq\frac{2M}{\text{min}(\delta_{1,\text{min}},\delta_{2,\text{min}})^{5/2}}\|\mathbf{D}_{1}-\mathbf{D}_{2}\|$
$\leq\frac{2M}{\text{min}(\delta_{1,\text{min}},\delta_{2,\text{min}})^{5/2}}\|\mathbf{A}_{1}\mathbf{1}_{p}-\mathbf{A}_{2}\mathbf{1}_{p}\|$
$\leq\frac{2Mp}{\text{min}(\delta_{1,\text{min}},\delta_{2,\text{min}})^{5/2}}\|\mathbf{A}_{1}-\mathbf{A}_{2}\|,$
where $M$ is the maximum norm of the adjacency matrix in the graph dataset, and $\mathbf{1}_{p}\in\mathbb{R}^{p}$ is the vector of ones.
Putting both inequalities together, we can come up with an upperbound for $\ell_{\mathrm{GCN}}$ that depends on the minimum degree in the dataset.
Appendix E Configuration Models
@@
In this section, we present a novel adaptation of configuration models as a graph data augmentation technique for GNN. Configuration models (Newman, 2013) enable the generation of randomized graphs that maintain the original degree distribution. We can, therefore, leverage this strategy to improve the generalization of GNNs. Below, we present the steps involved in our approach to using Configuration Models for Graph Data Augmentation:
@@
    Stub creation: Using a Bernoulli distribution with parameter $r\in[0,1]$, we randomly select a subset of candidate edges and break them to create stubs (half-edges).
@@
Table 4 shows the performance of this approach on the two GNN backbones, GCN and GIN.
Table 4: Classification accuracy ($\pm$ std) on different benchmark graph classification datasets for the data augmentation baselines based on the GIN backbone. The higher the accuracy (in %) the better the model.
Model IMDB-BIN IMDB-MUL MUTAG PROTEINS DD
Config Models w/ GCN $71.70{\scriptstyle\pm 3.16}$ $48.40{\scriptstyle\pm 3.88}$ $74.97{\scriptstyle\pm 6.77}$ $70.08{\scriptstyle\pm 4.93}$ $69.01{\scriptstyle\pm 3.44}$
Config Models w/ GIN $71.70{\scriptstyle\pm 4.24}$ $49.00{\scriptstyle\pm 3.44}$ $81.43{\scriptstyle\pm 10.05}$ $68.34{\scriptstyle\pm 5.30}$ $71.61{\scriptstyle\pm 5.93}$
As noticed, the configuration model-based graph augmentation method performs competitively with the baselines and even outperforms them in certain cases. This underscores the importance of Theorem 3.1. When compared to our approach GRATIN, the latter gives better results across different datasets and GNN backbones. This difference is primarily due to the configuration model based approach being model-agnostic, whereas GRATIN leverages the model’s weights and architecture, as explained in Section 3.4 and supported by Theorem 3.4.
Appendix F Ablation Study
To provide additional comparison and motivate the use of GMMs with the EM algorithm within GRATIN, we expanded our evaluation to include additional methods for modeling the distribution of the graph representations. Specifically, the comparison includes:
@@ Appendix F Ablation Study
Table 5: Ablation study on the density estimation scheme for learned GCN representations in GRATIN.
Model IMDB-BIN IMDB-MUL MUTAG PROTEINS DD
GMM w/ EM $\mathbf{71.00{\scriptstyle\pm 4.40}}$ $\mathbf{49.82{\scriptstyle\pm 4.26}}$ $\mathbf{76.05{\scriptstyle\pm 6.47}}$ $\mathbf{70.97{\scriptstyle\pm 5.07}}$ $\mathbf{71.90{\scriptstyle\pm 2.81}}$
GMM w/ VBI $\mathbf{71.00{\scriptstyle\pm 4.21}}$ $49.53{\scriptstyle\pm 4.26}$ $\mathbf{76.05{\scriptstyle\pm 6.47}}$ $\mathbf{70.97{\scriptstyle\pm 4.52}}$ $71.64{\scriptstyle\pm 2.90}$
KDE $55.90{\scriptstyle\pm 10.29}$ $39.53{\scriptstyle\pm 2.87}$ $66.64{\scriptstyle\pm 6.79}$ $59.56{\scriptstyle\pm 2.62}$ $58.66{\scriptstyle\pm 3.97}$
Copula $69.80{\scriptstyle\pm 4.04}$ $47.13{\scriptstyle\pm 3.45}$ $74.44{\scriptstyle\pm 6.26}$ $65.04{\scriptstyle\pm 3.37}$ $65.70{\scriptstyle\pm 3.04}$
GAN $70.60{\scriptstyle\pm 3.41}$ $48.80{\scriptstyle\pm 5.51}$ $75.52{\scriptstyle\pm 4.96}$ $69.98{\scriptstyle\pm 5.46}$ $66.26{\scriptstyle\pm 3.72}$
Table 6: Ablation study on the density estimation scheme for learned GIN representations in GRATIN.
Model IMDB-BIN IMDB-MUL MUTAG PROTEINS DD
GMM w/ EM $\mathbf{71.70{\scriptstyle\pm 4.24}}$ $\mathbf{49.20{\scriptstyle\pm 2.06}}$ $\mathbf{88.83{\scriptstyle\pm 5.02}}$ $\mathbf{71.33{\scriptstyle\pm 5.04}}$ $\mathbf{68.61{\scriptstyle\pm 4.62}}$
GMM w/ VBI $71.40{\scriptstyle\pm 2.65}$ $47.80{\scriptstyle\pm 2.22}$ $88.30{\scriptstyle\pm 5.19}$ $70.25{\scriptstyle\pm 4.65}$ $67.82{\scriptstyle\pm 4.96}$
KDE $69.10{\scriptstyle\pm 3.93}$ $41.46{\scriptstyle\pm 3.02}$ $77.60{\scriptstyle\pm 6.83}$ $60.37{\scriptstyle\pm 3.04}$ $67.48{\scriptstyle\pm 6.18}$
Copula $70.60{\scriptstyle\pm 2.61}$ $47.60{\scriptstyle\pm 2.29}$ $88.30{\scriptstyle\pm 5.19}$ $70.16{\scriptstyle\pm 4.55}$ $67.91{\scriptstyle\pm 4.90}$
GAN $70.50{\scriptstyle\pm 3.80}$ $48.40{\scriptstyle\pm 1.71}$ $\mathbf{88.83{\scriptstyle\pm 5.02}}$ $\mathbf{71.33{\scriptstyle\pm 5.55}}$ $67.74{\scriptstyle\pm 4.82}$
@@
Appendix G Training and Augmentation time
We compare the data augmentation times of our approach and the baselines in Table 7. In addition to outperforming the baselines on most datasets, our approach offers an advantage in terms of time complexity. The training time of baseline models varies depending on the augmentation strategy used, specifically, whether it involves pairs or individual graphs. Even in cases where a graph augmentation has a low computational cost for some baselines, training can still be time-consuming as multiple augmented graphs are required to achieve satisfactory test accuracy. For instance, methods like DropEdge, DropNode, and SubMix, while computationally simple, require generating multiple augmented samples at each epoch, thereby increasing the overall training time. Following the framework of (Yoo et al., 2022), we must sample several augmented graphs for each training graph at every epoch to achieve optimal results. In contrast, GRATIN  introduces a more efficient approach by generating only one augmented graph per training instance, which is reused across all epochs. This design ensures a balance between computational efficiency and augmentation effectiveness, reducing the overall training burden while maintaining strong performance. The only baseline that is more time-efficient than our approach is GeoMix; however, our method consistently outperforms GeoMix across all settings, as shown in Tables 1 and 2.
@@ Appendix G Training and Augmentation time
Time Model IMDB-BIN MUTAG DD